# Mixed-State Measurement-Induced Phase Transitions in Imaginary-Time Dynamics

### v2

@@ V.2 Stationary state
Figure 8: When setting $(\tau,g,p)=(1,3.5,0.5)$ for the MDITE with the 2D CDHM: (a) Convergence of the squared magnetization $\langle m^{2}\rangle$ with increasing circuit depth for various system sizes $L$. (b) Convergence of the Binder ratio $R_{2,z}$ with increasing circuit depth for various system sizes $L$.
V.3 Mixed-state phase transitions and analysis
The parameter space of the stationary state in this model is characterized by $(\tau,g,p)$. Compared with the TFIM Hamiltonian, the CDHM Hamiltonian not only provides a two-dimensional example but also exhibits a global continuous $SU(2)$ symmetry. Interestingly, even at $g=0$, the ITE operator contains off-diagonal components $(S^{+}_{i}S^{-}_{j}+S^{-}_{i}S^{+}_{j})/2$. Increasing $g$ introduces both diagonal and off-diagonal terms, in sharp contrast to the TFIM case where tuning the Hamiltonian parameter $h$ only modifies the off-diagonal part.
Following the same strategy as in the TFIM case, we fix two of the parameters in $\{\tau,g,p\}$ and vary the third to search for mixed-state phase transitions. Taking the example of tuning $p$ at fixed $(\tau,g)=(1,3.5)$ in Fig. 9(a) and (b), we observe a clear crossing of the Binder ratio $R_{2,z}$ at $p_{c}\approx 0.354$, with $R_{2,z}\to 3$ for $p<p_{c}$ and $R_{2,z}\to 1$ for $p>p_{c}$. If $p=0$, the system is in the dimerized phase which has no order in the staggered magnetization. This indicates a mixed-state phase transition from a disordered phase to an AFM Ising-like ordered phase along the $z$-axis, driven by increasing the measurement strength. Although the Binder ratio exhibits Ising-like characteristic, similar to the case of TFIM, the extracted critical exponents in the case of 2D CDHM do not correspond to any known universality class including the 3D Ising universality class for a closed system, as shown in Table 4. Similar phenomena can also be observed when tuning $g$ and $\tau$, as shown in Fig. 9(c-f) and Tables 5 and 6.
Importantly, the critical exponents $\nu$ and $\beta$ also differ as we tune $p$, $\tau$, or $g$, but the ratio $\beta/\nu\approx 0.9$ across all cases remains consistent within error bars in our simulations, as shown in Tables 4, 5, and 6. Therefore, we also argue that these mixed-state phase transitions in the 2D CDHM belong to a single universality class when $p>0$, despite the differences in $\nu$ and $\beta$.
To interpret these results, note that any nonzero measurement rate $p$ explicitly breaks the $SU(2)$ symmetry to a residual $U(1)\rtimes\mathbb{Z}_{2}$, favoring the $z$-axis. According to the Mermin–Wagner theorem [Mermin1966, Hohenberg1967, Sondhi1997Continuous], a continuous $U(1)$ symmetry cannot undergo spontaneous symmetry breaking in two dimensions for any finite imaginary-time extent. Consequently, the transition observed here can only break the discrete symmetry $\mathbb{Z}_{2}$ induced by the projective measurements. Although increasing $g$ in the Hamiltonian introduces both diagonal and off-diagonal terms, only the diagonal components are compatible with the projective measurements along the $z$-axis, which correspond to the AFM Ising interactions. In Appendix. D, we present the results demonstrating the absence of off-diagonal long-range order in this model.
When $g>g_{c,\mathrm{GS}}$, and the imaginary-time steps are short ($\tau g\ll 1$), a high measurement rate $p$ suppresses the off-diagonal coherence generated by the ITE operator, leading to dynamics dominated by the diagonal AFM Ising terms. This drives the system into an ordered phase with nonzero staggered magnetization along the $z$-axis. In contrast, if $\tau g\gg 1$, the off-diagonal coherence builds up rapidly, and the final stationary state remains disordered even at $p=1$. Therefore, for projective measurements along the $z$-axis, the MDITE with the 2D CDHM effectively generalizes the physics of the 1D TFIM to two dimensions.
Since the CDHM can also be simulated using cluster updates within QMC, the emergence of long-range orders induced by measurements can naturally be understood qualitatively through the cluster formation mechanism, as discussed in Appendix. C. We emphasize that it would be interesting to explore other types of decoherence channels that preserve other symmetries, such as the $SU(2)$ symmetry, which we leave for future work. In such cases, more advanced update algorithms beyond conventional cluster methods for QMC stimulations would be necessary, potentially revealing qualitatively different types of phase transitions. Physically, the enhancement of long-range order induced by measurement can be interpreted as a suppression of quantum fluctuations—i.e., the off-diagonal elements of the density matrix—caused by the diagonal measurement employed in our setup.
The schematic critical surface associated with this model is shown in Fig. 9(g) and (h), which provides a qualitative view to understand the phase transition driven by $p$, $g$ and $\tau$. This critical surface also shows that large $g$ requires small evolution time, while long evolution time requires weak coupling $g$. This inverse relationship between $g$ and $\tau$ closely resembles that of the 1D TFIM.
@@ V.3 Mixed-state phase transitions and analysis
(g) and (h): Critical surfaces viewed from different angles. To facilitate the distinction of critical $p$ corresponding to different parameters on the critical surface, the figure employs a colorbar to map $p$ to different colors.
$L_{\min}$ $p_{c}$ $\nu$ $\beta$ $\beta/\nu$
$24$ 0.3537(4) 0.714(7) 0.650(11) 0.910(7)
$28$ 0.3537(4) 0.716(6) 0.651(11) 0.909(8)
$32$ 0.3538(5) 0.717(6) 0.650(12) 0.906(9)
Table 4: We set $(\tau,g)=(3.5,1)$ to extract the critical exponent of $\langle|m|\rangle$ in the MIPT by using system size $L=24,28,32,40,48$. To test the stability of the extracted exponent, we gradually discard the small system size to perform the data collapse. $L_{\min}$ is the smallest system size used in the data collapse from 24 to 32. The resulting stable values are $p_{c}\approx 0.354$, $\nu\approx 0.714$, $\beta\approx 0.65$, and $\beta/\nu\approx 0.90$.
$L_{\min}$ $\tau_{c}$ $\nu$ $\beta$ $\beta/\nu$
$24$ 0.4679(5) 0.799(6) 0.718(9) 0.898(5)
$28$ 0.4679(6) 0.801(6) 0.721(10) 0.900(7)
$32$ 0.4678(7) 0.801(6) 0.720(12) 0.899(9)
Table 5: We set $(g,p)=(3,0.1)$ to extract the critical exponent of $\langle|m|\rangle$ in the MIPT by using system size $L=24,28,32,40,48$. To test the stability of the extracted exponent, we gradually discard the small system size to perform the data collapse. $L_{\min}$ is the smallest system size used in the data collapse from 24 to 32. The resulting stable values are $\tau_{c}\approx 0.4679$, $\nu\approx 0.8$, $\beta\approx 0.72$, and $\beta/\nu\approx 0.90$.
$L_{\min}$ $g_{c}$ $\nu$ $\beta$ $\beta/\nu$
$24$ 2.806(3) 0.736(18) 0.654(26) 0.89(2)
$28$ 2.806(2) 0.739(17) 0.653(26) 0.88(2)
$32$ 2.806(2) 0.741(15) 0.650(26) 0.88(2)
Table 6: We set $(\tau,p)=(2,0.3)$ to extract the critical exponent of $\langle|m|\rangle$ in the MIPT by using system size $L=24,28,32,40,48$. To test the stability of the extracted exponent, we gradually discard the small system size to perform the data collapse. $L_{\min}$ is the smallest system size used in the data collapse from 24 to 32. The resulting stable values are $g_{c}\approx 2.806$, $\nu\approx 0.74$, $\beta\approx 0.65$, and $\beta/\nu\approx 0.90$.
VI Conclusion and discussions
@@
In this work, we have introduced MDITE as a versatile framework for studying novel mixed-state phases and phase transitions in open quantum systems. By combining ITE with probabilistic projective measurements, MDITE enables the exploration of fully non-unitary dynamics and their stationary states, providing both conceptual insight and numerical accessibility for studying high-dimensional open quantum systems.
@@
As illustrative examples, we applied MDITE to the 1D TFIM and the 2D CDHM, uncovering distinct mixed-state phases and associated phase transitions in the stationary states of the protocol. Remarkably, the Ising-like universalities observed in both 1D and 2D models are characterized by the ratio of critical exponents $\beta/\nu$, rather than the individual exponents. To the best of our knowledge, these represent new universality classes in open quantum systems. We provided a visualized understanding of these criticalities and further showed that the transitions can be interpreted in terms of cluster formation within the QMC simulations. These results demonstrate that the MDITE is an excellent framework for exploring nontrivial mixed states, yielding fresh insights into the physics of open quantum systems.
@@
At the methodological level, we introduced a diagrammatic representation of the MDITE process, which facilitates QMC and other numerical simulations on a broad class of many-body Hamiltonians, including those exhibiting spontaneous symmetry breaking, topological phase transitions, and topologically ordered phases. While our analysis focused on magnetization and the Binder ratio, the computation of information-theoretic quantities—such as entanglement entropy [hasting2010renyi, stephan2012renyi, melko2010renyi, jonathan2020renyi, Zhao2022renyi, ding2024reweight, wangz2025reweight, tarabunga2025bell, yan2023unlocking, li2024relevant, mao2025sampling, wang2025extracting] and Rényi negativity (including entanglement witnesses based on partitial transpose moments) [ding2025tracking, Wukh2020negativity, Alba2013negativity, Wangfh2025negativity, chungchiamin2014negativity, yu2021pt, elben2020pt, Neven2021pt, tarabunga2025pt]—is straightforward in combination with recent QMC techniques.
@@
Our work opens several promising avenues for future research. One direction is to further characterize the universality classes associated with the criticalities observed in our simulations. Moreover, our QMC approach allow systematic investigations of dynamical properties and nonequilibrium phenomena of the MDITE process. Another is to extend the framework to more complex Hamiltonians and general decoherence channels beyond projective measurements, thereby accessing a broader landscape of many-body phenomena. In particular, long-range entanglement or topological order in mixed states [lu2023charac, lu2023longrange, Fanruihua2024diagnostics, Lu2024disentangling, sang2025markov, lessa2025higherformanomaly], as well as the strong-to-weak spontaneous symmetry-breaking [Lee2023decoh_criticality, Lessa2025swssb, sala2024swssb, gu2024swssb, huang2025swssb, zack2025renyi1corr, guo2025swssb, guoyc2025swssb, guoyc2025ldpo, zhang2025swssb, zpf2025swssb, Liu2025swssb] can be explored within the MDITE framework.
@@
On the experimental side, MDITE could be implemented in platforms such as Rydberg atom arrays [Wu2021rydbergreview, Evered2023rydberg], superconducting qubits [Wendin2017superconduct], or trapped-ion systems [haffner2008trapped]. Such realizations would enable direct observation of the mixed-state criticalities predicted here, bridging theory and experiment.

### v3

@@ V.2 Effective Hamiltonian in the combined limit
For the example of the 1D TFIM (10), the effective Hamiltonian takes the form
${H}_{\text{eff}}=$ $-\frac{1}{2}\sum_{\langle ij\rangle}(Z_{i}^{L}Z_{j}^{L}+Z_{i}^{R}Z_{j}^{R})-\frac{h}{2}\sum_{i}(X_{i}^{L}+X_{i}^{R})$ (23)
$-\frac{p}{2\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}.$
When $p/\tau=0$, the two copies are decoupled and the $H_{\text{eff}}$ has a $\mathbb{Z}_{2}^{L}\times\mathbb{Z}_{2}^{R}$ symmetry, where $\mathbb{Z}_{2}^{L(R)}$ corresponds to independent spin-flip symmetry in the left (right) copy, generated by $X^{L(R)}\equiv\prod_{i}X_{i}^{L(R)}$.
In the combined limit $\tau\sim\tilde{\tau}\sim p$, the inter-copy coupling is finite, and only the diagonal spin-flip transformation preserves the Hamiltonian. The symmetry is therefore reduced to the diagonal subgroup $\mathbb{Z}_{2}^{\text{diag}}=\{1,X^{L}X^{R}\}\cong\mathbb{Z}_{2}$, and the effective Hamiltonian (23) is equivalent to a standard TFIM defined on a two-leg ladder geometry, as illustrated in Fig. 8(a).
Figure 8: (a) Ladder geometry of $H_{\text{eff}}$ in Eq. (23) for the case where $H$ is the 1D TFIM. The two legs correspond to the left ($L$) and right ($R$) copies of the chain (two rings under periodic boundary conditions). The interlayer Ising couplings along the rungs have strength $p/2\tau$. (b) Ground-state phase diagram of $H_{\text{eff}}$ obtained from QMC simulations.
Fig. 8(b) shows the ground-state phase diagram of $H_{\text{eff}}$ in Eq. (23). The critical point at $h=1$ lies at $p/\tau=0$. Note that the combined limit requires $p/\tau\sim O(1)$, and this corresponds to the region $h>1$. Thus, sufficiently large $p/\tau$ can drive a transition from the disordered phase to the ordered FM phase when $h>1$ in the doubled space.
V.3 MIPT in the combined limit
Note that in the effective Hamiltonian (23), both the interlayer Ising coupling (from measurements) and the intralayer Ising coupling (from the Ising interaction in the original Hamiltonian) favor FM ordering in the same diagonal $\mathbb{Z}_{2}$ channel, while the transverse field $h$ enhances quantum fluctuations and tends to disorder the system.
The key to understanding the mixed-state MIPT in the combined limit is the dual role of the $Z$-basis: it is both the basis in which projective measurements are performed and the basis in which the ordered phase of the Hamiltonian is defined. When $p/\tau$ is sufficiently large compared to hh—equivalently, when the decoherence strength $p$ dominates over the coherence scale $\tau h$ via the ITE operator $e^{-\frac{\tau}{2}H}$—the system develops FM long-range order for the steady state in the doubled space.
This ordered phase is characterized by the local order parameter $\langle\!\langle\rho_{\infty}|Z_{i}^{L(R)}|\rho_{\infty}\rangle\!\rangle$ (or one can equivalently consider the magnetization averaged over all sites), where $\left|\rho_{\infty}\right\rangle\!\rangle$ denotes the ground state of $H_{\text{eff}}$. In the original Hilbert space, the Choi mapping relates this quantity to the nonlinear quantity $\mathrm{Tr}(\rho_{\infty}^{2}Z_{i})$, with $\rho_{\infty}$ denoting the steady state of the MDITE. Though the parameter region where $\mathrm{Tr}(\rho_{\infty}^{2}Z_{i})\neq 0$ needs not coincide exactly with that where the linear order parameter $\mathrm{Tr}(\rho_{\infty}Z_{i})$ is nonzero, both capture the SSB physics. The resulting mixed-state MIPT for the steady state of the MDITE protocol in the combined limit is therefore expected to be qualitatively identical and within the same 2D classical Ising universality class.
We emphasize that the results in the combined limit are qualitatively consistent with the numerical results in Sec. IV.3 away from the continuous limits. The mechanism of these mixed-state phase transitions in Sec. IV.3 could therefore be understood as follows.
V.4 MIPT away from the continuous limits
In the regime $h>1$ with relatively short imaginary-time length ($\tau h\lesssim p$), frequent projective measurements at large $p$ interrupt the buildup of quantum coherence. The transverse field, which generates quantum fluctuations in the $X$ basis, is repeatedly suppressed before correlations can develop, as the state is frequently projected onto the $Z$ basis. With off-diagonal coherence effectively suppressed, the dynamics become dominated by the diagonal Ising terms, steering the system toward a FM phase analogous to that of the classical Ising model.
In contrast, when the transverse field is sufficiently strong or the evolution time is sufficiently long ($\tau h\gtrsim p$), quantum coherence develops rapidly within each imaginary-time interval. Coherence is established faster than it can be suppressed by measurements, preventing the formation of FM order even at $p=1$. Consequently, no phase transition occurs in this regime upon tuning the measurement rate $p$. One may also expect that when $h\leq 1$, no phase transition characterized by spin magnetization occurs upon tuning $p$, which has also been observed in our numerical observations.
Since the competition between quantum coherence and measurement-induced decoherence can establishe a critical balance, tuning the system parameters, including $p$, $\tau$, and $h$, toward this balance drives a mixed-state phase transition separating the ordered and disordered phases, as shown in Sec. IV.4 In this sense, transitions induced by varying $p$, $\tau$, or $h$ originate from the same underlying mechanism. The difference lies only in which parameter controls the competition.
Furthermore, while the combined limit admits a description in terms of the effective Hamiltonian $H_{\mathrm{eff}}$ and exhibits critical behavior consistent with the Ising universality class, the numerical results in the intermediate regime (away from the three limits) show critical exponents that do not match those of any known universality class. In particular, the extracted exponents appear to vary continuously as the parameters $h$, $\tau$, and $p$ are changed. These observations suggest the possibility that the mixed-state phase transitions in the intermediate regime belong to some distinct universality class that has not been previously identified.
Physically, the nontrivial mixed-state phase transitions should arise from the intrinsically nonequilibrium nature of the problem. Based on the BCH expansion, one can also define an effective Hamiltonian $H_{\text{eff}}$ through
$e^{-\tau\tilde{H}}e^{\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}}\equiv e^{-\tau H_{\text{eff}}}.$ (24)
However, In general, the resulting effective Hamiltonian $H_{\text{eff}}$ could become highly nonlocal due to the infinite series of nested commutators, resulting critical behavior therefore fall outside the standard equilibrium classification.
In addition, the emergence of long-range order with an increasing measurement rate can also be understood intuitively from the perspective of QMC simulations: a higher measurement rate naturally generates larger clusters that connect a broader range of sites in each QMC update. This provides a direct and intuitive picture of how measurements drive ordering in the system of TFIM, highlighting the use of the diagrammatic representations in Sec. III and the advantage of QMC in revealing the underlying mechanism. Further details are provided in Appendix. E.
Figure 9: Away from the continuous limits, critical surfaces of the steady state of the MDITE protocol viewed from different angles [(a) and (b)]. To facilitate the distinction of critical $p$ corresponding to different parameters on the critical surface, the figure employs a colorbar to map $p$ to different colors.
V.5 Critical surface away from the continuous limits
To visualize the phase transitions of the MDITE steady states, we construct a schematic phase diagram based on approximate transition points obtained from Binder ratio crossings in several system sizes, as shown in Fig. 9. Importantly, Fig. 9(b) shows that the critical surface shifts toward larger $\tau$ as $h$ approaches $1$. This trend is expected: longer evolution times require weaker transverse fields to allow FM order to emerge under strong measurements. Consequently, in the limit $h\to 1$, the critical surface is expected to extend infinitely along the $\tau$-axis, permitting a transition at any $\tau$ by tuning $p$. A similar reasoning applies in the opposite limit, $h\to\infty$ and $\tau\to 0$.
VI Two-dimensional model with continuous symmetry
VI.1 Model and observables
As the second example, we consider the 2D $S=1/2$ columnar dimerized Heisenberg model (CDHM) on a $L\times L$ square lattice [59]. The Hamiltonian is given by
where $\mathbf{S}_{i}$ is the spin operator at site $i$, and $\langle ij\rangle_{1}$ and $\langle ij\rangle_{g}$ denote two different sets of nearest-neighbor pairs, as shown in Fig. 10. We also consider the local $Z$-basis as the computational basis, and the periodic boundary condition is chosen. The model has a global $SU(2)$ symmetry and the ground state exhibits a quantum phase transition at $g_{c,\mathrm{GS}}\approx 1.90951$ [81, 59, 55], which belongs to the 3D $O(3)$ universality class. The critical point separates a Néel ordered phase ($g<g_{c,\mathrm{GS}}$) from a dimerized phase ($g>g_{c,\mathrm{GS}}$).
Due to the antiferromagnetic (AFM) nature of the Heisenberg interactions, we consider the staggered magnetization $\langle m_{z}\rangle=\Big\langle\sum_{i}(-1)^{x_{i}+y_{i}}Z_{i}/N\Big\rangle$ as the order parameter, where $x_{i}$ and $y_{i}$ denote the integer coordinates of site $i$ in the lattice. As in the TFIM, we compute the absolute magnetization $\langle|m_{z}|\rangle$, the second and fourth moments of the magnetization, and the Binder ratio $R_{2,z}$. Furthermore, for the ground state of the CDHM, the Binder ratio associated with the $O(3)$ phase transition approaches $R_{2,z}\to 9/5$ in the Néel phase and $R_{2,z}\to 3$ in the dimerized phase [78, 105, 17, 55].
VI.2 Steady state
As in the 1D TFIM example, the MDITE protocol in this case also drives the system toward a steady state in the limit of large number of time steps $n_{d}$. Fig. 11 illustrates this convergence for representative parameters $(\tau,g,p)=(1,3.5,0.5)$, where both the staggered magnetization $\langle|m_{z}|\rangle$ and the Binder ratio $R_{2,z}$ quickly approach their stationary values. In practice, we again find that $n_{d}=2L/\tau$ is sufficient to reach the steady state, and this choice is used in all subsequent QMC simulations for the CDHM. As discussed later, the dynamical exponent $z$ for the mixed-state phase transitions in this example is also around one.
@@ VI.2 Steady state
Figure 11: When setting $(\tau,g,p)=(1,3.5,0.5)$ for the MDITE with the 2D CDHM: (a) Convergence of the squared magnetization $\langle m^{2}\rangle$ with increasing the number of time steps $n_{d}$ for various system sizes $L$. (b) Convergence of the Binder ratio $R_{2,z}$ with increasing $n_{d}$ for various system sizes $L$.
VI.3 Mixed-state phase transitions
@@
(g) and (h): Critical surfaces viewed from different angles.
The parameter space of the steady state in this model is spanned by $(\tau,g,p)$. Compared with the TFIM Hamiltonian, the CDHM Hamiltonian not only provides a 2D example but also exhibits a global continuous $SU(2)$ symmetry. Interestingly, even at $g=0$, the ITE operator contains off-diagonal components $(S^{+}_{i}S^{-}_{j}+S^{-}_{i}S^{+}_{j})/2$. Increasing $g$ introduces both diagonal and off-diagonal terms, in sharp contrast to the TFIM case where tuning the Hamiltonian parameter $h$ only modifies the off-diagonal part. Following the same strategy as in the TFIM case, we fix two of the parameters in $\{\tau,g,p\}$ and vary the third to search for mixed-state phase transitions.
Taking the example of tuning $p$ at fixed $(\tau,g)=(1,3.5)$ in Fig. 12(a) and (b), we observe a clear crossing of the Binder ratio $R_{2,z}$ at $p_{c}\approx 0.354$, with $R_{2,z}\to 3$ for $p<p_{c}$ and $R_{2,z}\to 1$ for $p>p_{c}$. If $p=0$, the system is in the dimerized phase which has no order in the staggered magnetization. This indicates a mixed-state phase transition from a disordered phase to an AFM Ising-like ordered phase along the $z$-axis, driven by increasing the measurement strength.
Although the Binder ratio exhibits Ising-like characteristic, similar to the case of TFIM, the extracted critical exponents in the case of 2D CDHM again do not correspond to any known universality class including the 3D Ising universality class for a closed system, as shown in Table 3. Similar phenomena can also be observed when tuning $g$ and $\tau$, as shown in Fig. 12(c-f) and Table 3. Importantly, the critical exponents $\nu$ and $\beta$ also change as we tune $p$, $\tau$, or $g$.
Parameters $L_{\min}$ Critical point $\nu$ $\beta$
$(\tau,g)=(3.5,1)$ 24 $p_{c}=0.3537(4)$ 0.714(7) 0.650(11)
28 $p_{c}=0.3537(4)$ 0.716(6) 0.651(11)
32 $p_{c}=0.3538(5)$ 0.717(6) 0.650(12)
$(g,p)=(3,0.1)$ 24 $\tau_{c}=0.4679(5)$ 0.799(6) 0.718(9)
28 $\tau_{c}=0.4679(6)$ 0.801(6) 0.721(10)
32 $\tau_{c}=0.4678(7)$ 0.801(6) 0.720(12)
$(\tau,p)=(2,0.3)$ 24 $g_{c}=2.806(3)$ 0.736(18) 0.654(26)
28 $g_{c}=2.806(2)$ 0.739(17) 0.653(26)
32 $g_{c}=2.806(2)$ 0.741(15) 0.650(26)
Table 3: Finite-size stability analysis of the critical exponents obtained from data collapse of $\langle|m|\rangle$ in the CDHM. System sizes are $L=24,28,32,40,48$. Small system sizes are progressively discarded to test the stability of the extracted critical exponents.
We also extract dynamical exponents $z\approx 1$ for all three quantum critical points. These results are consistent with approximately isotropic space-time scaling and are compatible with an emergent Lorentz-invariant description. Further details are provided in Appendix G.
VI.4 Analysis of the phase transitions
To further interpret these results, note that any nonzero measurement rate $p$ explicitly breaks the (weak) $SU(2)$ symmetry to a residual (weak) $U(1)\rtimes\mathbb{Z}_{2}$ symmetry, favoring the $z$-axis.
Although increasing $g$ in the Hamiltonian introduces both diagonal and off-diagonal terms, only the diagonal components are compatible with the projective measurements along the $z$-axis, which correspond to the AFM Ising interactions. In Appendix. H, we also present the results demonstrating the absence of off-diagonal long-range order in this model.
To understand this kind of phase transitions, one can similarly construct an effective two-copy Hamiltonian in the combined limit $\tau\to 0$, $p\to 0$, with $\tau\sim p$, as that in Sec. V. In this limit, the resulting effective model corresponds to a bilayer geometry consisting of two CDHM layers coupled by interlayer Ising interactions. When each individual layer is in the dimerized disordered phase ($g>g_{c,\text{GS}}$), tuning the measurement rate $p$ (equivalently, the interlayer coupling strength) can drive Ising phase transitions associated with the 3D classical Ising universality class.
Away from the continuous limits, when $g>g_{c,\mathrm{GS}}$, and the imaginary-time length is short ($\tau g\lesssim p$), a high measurement rate $p$ suppresses the off-diagonal coherence generated by the ITE operator, leading to dynamics dominated by the diagonal AFM Ising terms. This drives the system into an ordered phase with nonzero staggered magnetization along the $z$-axis. In contrast, if $\tau g\gtrsim p$, the off-diagonal coherence builds up rapidly, and the final steady state remains disordered even at $p=1$.
Therefore, for projective measurements along the $z$-axis, the MDITE with the 2D CDHM generalizes the physics of the 1D TFIM discussed in Secs. IV and V to 2D. The associated phase transitions also appear to lie beyond conventional universality classes.
Since the CDHM can also be simulated using cluster updates within QMC, the emergence of long-range orders induced by measurements can also be understood through the cluster formation in Appendix. E. We emphasize that it would be interesting to explore other types of decoherence channels that preserve other symmetries, such as the $SU(2)$ symmetry, which we leave for future work. In such cases, more advanced update algorithms beyond conventional cluster methods for QMC simulations would be necessary, potentially revealing qualitatively different types of phase transitions.
The schematic critical surface associated with this model is shown in Fig. 12(g) and (h). This critical surface also shows that large $g$ requires small evolution time, while long evolution time requires weak coupling $g$. This inverse relationship between $g$ and $\tau$ closely resembles that of the 1D TFIM.
Moreover, we note that since the state still has a $U(1)$ symmetry, off-diagonal long-range order may emerge if the imaginary-time step length in the MDITE protocol is allowed to scale with the system size, i.e., $\tau\propto L^{\alpha}$ for a sufficiently large $\alpha$. In this regime and for $g<g_{c,\text{GS}}$, the final step of the MDITE protocol (measurement followed by ITE) effectively reduces to a conventional ITE procedure for preparing the ground state or low-energy excited state of the Hamiltonian, with the initial state to be $\mathcal{E}_{p}(\rho_{n_{d}-1})$. This regime represents an interesting extension of the present framework and may be explored in future studies.
VII Conclusion and discussions
@@
In this work, we have introduced MDITE as a versatile framework for studying novel mixed-state phases and phase transitions in open quantum systems. By combining ITE with probabilistic projective measurements, MDITE enables the exploration of fully non-unitary dynamics and their steady states, providing both conceptual insight and numerical accessibility for studying high-dimensional open quantum systems.
@@
As illustrative examples, we applied MDITE to the 1D TFIM and the 2D CDHM, uncovering distinct mixed-state phases and associated phase transitions in the steady states of the protocol. Remarkably, the Ising-like critical behavior observed in both 1D and 2D models may, to the best of our knowledge, represent new universality classes of mixed-state phase transitions in open quantum systems. We also developed a theoretical description in the important combined limit in Sec. V, which yields results qualitatively consistent with the numerical observations. In addition, the transitions admit a natural interpretation in terms of cluster formation within the QMC simulations, providing an intuitive picture of the emergent ordering in the mixed states.
@@
At the methodological level, we introduced a diagrammatic representation of the MDITE process, which facilitates QMC and other numerical simulations on a broad class of many-body Hamiltonians, including those exhibiting SSB, topological phase transitions, and topologically ordered phases. While our analysis focused on magnetization and the Binder ratio, the computation of information-theoretic quantities—such as entanglement entropy [35, 38, 61, 15, 127, 18, 108, 96, 114, 46, 57, 107] and Rényi negativity (including entanglement witnesses based on partitial transpose moments) [19, 112, 1, 104, 14, 120, 22, 67, 97]—is straightforward in combination with recent QMC techniques.
Our work opens several promising directions for future research. First, establishing the precise universality classes of the mixed-state phase transitions observed here remains an important open problem. In particular, it would be valuable to understand how these critical points, which appear to go beyond conventional Ising universality, connect to standard Ising critical points in the three continuous limits in Sec. V.
@@
Second, it would be interesting to extend the MDITE framework to more complex Hamiltonians and more general decoherence channels beyond projective measurements, thereby accessing a broader range of many-body phenomena. In particular, long-range entanglement or topological order in mixed states [52, 53, 24, 54, 85, 45], as well as strong-to-weak spontaneous symmetry breaking [43, 44, 76, 29, 37, 109, 33, 31, 32, 124, 91, 50, 126, 83], can be explored within the MDITE framework.
While we have characterized the dynamical critical exponent in the present work, it would be interesting to further explore more detailed dynamical properties of the protocol. In particular, a systematic study of relaxation dynamics and critical scaling toward the steady state may reveal richer structures and additional universal features beyond those captured by the stationary mixed states.
@@
On the experimental side, MDITE could be implemented in platforms such as Rydberg atom arrays [113, 23], superconducting qubits [110], or trapped-ion systems [34]. Such realizations would enable direct observation of the mixed-state criticalities predicted here, bridging theory and experiment. Therefore, the MDITE framework may provide a useful platform for designing quantum algorithms for preparing nontrivial mixed states, potentially enabling efficient simulations of open quantum many-body systems.